Alcan invested $60 million in a new packaging facility in North Carolina. If the plant is to be depreciated over 10 years with straight line depreciation, sales are to generate $43 million in revenues per year, operating and maintenance costs are $26 million per year, and there is no salvage value, what is the after-tax cash flow (in millions of $) from the year 8 of production for the facility? Assume an effective tax rate of 35%.
Answer #1

Initial investment=I=$60 million

Salvage=S=0

Useful life=n=10 years

According to straight line depreciation,

Depreciation per year=(I-S)/n=(60-0)/10=$6 million

Sales per year=S=$43 million

O&M expenses per year=O&M=$26

Before Tax cash flow at the end of year 8=BTCF=S-O&M=43-26=$17 million

Tax liability=(BTCF-D)*Tax Rate=(17-6)*35%=$3.85 million

After tax cash flow at the end of year 8=BTCF-Tax liability=17-3.85=$13.15 million